there are two answers posted for this question. can you confirm which one is the correct one and why? ---------------------- question: You are considering a new product launch. The project will cost $950,000 and have a four-year life with no salvage value. Assets for the project are depreciated straight line down to zero. Projected sales are 220 units per year. The price per unit is $18,400 while variable cost per unit is $14,900. Fixed costs per year are $320,000. The firm
Answer #1
answer 1: EBDIT= 220(18400-14900)-320000= $450000 Less dep: (950000/4)= 237500 EBT= 212500 EAT(65%)= 138125 Cashflow= 138125+237500= $375625 PV= 375625/1.12+(375625/1.12^2)+(375625/1.12^3)+(375625/1.12^4)= $1140904.35
This is the correct answer.
Answer #2
Hi, I believe the second answer was posted by me. However, it is not correct, because I did not took into account the amount of depreciation. Correct answer is as below: OCF = [(18,400